OpenStudy (anonymous):

find the slope of M(-9,16) N(-11,16) and the perpendicular slope thanks :)

OpenStudy (cwrw238):

slope of line joining M and N = 16-16/-11--9 = 0 - the line is parallel tp x-axis line perpendicular is parallel to y-axis

OpenStudy (anonymous):

@cwrw238 how did u get 0? when the denominator is still a whole number?

OpenStudy (cwrw238):

the numerator is 16-16 = 0 0 / -2 = 0 0 divided by any number which is not zero = 0

OpenStudy (anonymous):

o i get it now but to find the reciprocal inverse your have to flip it and then multiply by -1 how do you do the next step?

OpenStudy (anonymous):

@thechocoluver445 could u please help me?

OpenStudy (anonymous):

\[Recall\: that\: slope\: is\: \frac{y_2-y_1}{x_2-x_1}\]\[\And\: perpendicular\: slope\: is\: the\: opposite\: recpirocal: For\: example:\: If\: the\: slope\: is\: -\frac{2}{3}, then\: the\: opposite\: recpirocal\: is \frac{3}{2}\]

OpenStudy (anonymous):

yea but the slope is 0 and how can i reciprocal 0?

OpenStudy (anonymous):

If the slope is 0, that means it has to be a horizontal line:  Horizontal lines are perp. to vertical lines, so that means the slope is undefined (vertical lines have undefined slope)

OpenStudy (anonymous):

@thechocoluver445 okay so the perpendicular slope is undefined as well?

OpenStudy (anonymous):

The slope of the line is zero The perpendicular slope is undefined Yeah.

OpenStudy (anonymous):

thankyou so much :)

OpenStudy (anonymous):

yw :)